Multiply the equation by a power of ten to write an equivalent equation with integer coefficients.
2.5 x 0.7=4.6-1.3x
1.67+2.43x=3.29(x-5)

Answer by stanbon(48535)

2.5 x 0.7=4.6-1.3x
1.67+2.43x=3.29(x-5)
------------------------
Simplify 1st and multiply 2nd thru by 100 to get:
1.75 = 4.6 - 1.3x
167 + 243x = 329(x-5)
--------------------------
Multiply thru 1st by 100 and leave 2nd alone to get:
175 = 460 - 130
167 + 243x = 329(x-5)
=============================
Cheers,
Stan H.